# A $\mathbb{CP}^{2}$ SMEFT

### v2

@@ 1 Introduction
The Lagrangian has terms involving: (i) Gauge-covariant derivatives of the Higgs field, (ii) a Higgs potential, (iii) Gauge field tensors, (iv) Dirac (kinetic) terms for fermions, (v) Yukawa terms coupling the Higgs field and fermions. We will discuss these in turn, and then explore some phenomenological consequences, in particular, for the $W$-boson to $Z$-boson mass ratio, and for the Yukawa couplings of the Higgs boson. These will lead to constraints on the BSM mass parameter $M$.
@@ 2 $\mathbb{CP}^{2}$ Geometry
The complex projective plane $\mathbb{CP}^{2}$ is obtained from flat complex 3-space, $\mathbb{C}^{3}$, by quotienting out the action of scaling by a complex number. Geometrically, $\mathbb{CP}^{2}$ is the manifold of complex lines [^2] We use the mathematician’s notion of complex line $\mathbb{C}$ to denote what is more commonly called the complex plane. All lines considered here are lines through the origin of $\mathbb{C}^{3}$. in $\mathbb{C}^{3}$.
@@
Later, we will be using the projections of a generic vector $\Psi\in\mathbb{C}^{3}$, defined at location $Z\in\mathbb{C}^{3}$, on to the line through $Z$ and on to its complement, which is a copy of the tangent space to $\mathbb{CP}^{2}$. These projections are respectively
@@
Because these projections are unaffected by rescaling $Z$ by $\kappa$, they are defined at location $Z$ on $\mathbb{CP}^{2}$. More precisely, over $\mathbb{CP}^{2}$ there is its 1-dimensional tautological bundle, whose fibre at a point $Z\in\mathbb{CP}^{2}$ is the line in $\mathbb{C}^{3}$ through $Z$, and its 2-dimensional tangent bundle. The 3-dimensional, trivial $\mathbb{C}^{3}$ bundle over $\mathbb{CP}^{2}$ is the direct sum of these, so an element $\Psi$ in the trivial $\mathbb{C}^{3}$ bundle is the sum of its projections on to the tautological and tangent bundles.
@@ 3 Squashed 3-Spheres
whose angular part is the metric of the round 3-sphere of radius $r$. Each of these 3-spheres is a $U(2)$ orbit. and there is a unique $U(2)$ fixed point at $r=0$. Note that the ratio of the circle radius to the 2-sphere radius is a constant, independent of $r$.
@@
Here, the angular part is the metric of a squashed 3-sphere, a Hopf circle bundle over a round 2-sphere, where the ratio of the circle radius to the 2-sphere radius is $1/\sqrt{1+r^{2}}$, i.e. dependent on $r$. For us, the important observation is that for small positive $r$, the radius ratio is close to unity, so the squashed 3-sphere is approximately round. The consequence for our model is that the custodial $SO(4)$ symmetry – the symmetry of a round 3-sphere – will be approximately realised by the Higgs potential.
As $r\to\infty$, the radius ratio approaches zero, compatibly with the completion of $\mathbb{CP}^{2}$ at infinity being simply a round 2-sphere, which is a special 2-dimensional orbit of $U(2)$. It has radius $\frac{1}{2}$ and is at a finite distance from any other point in $\mathbb{CP}^{2}$.
@@ 4 Higgs Kinetic Term
where $\phi=(\phi_{1},\phi_{2})$ is a complex scalar doublet.333 All these multiplets, and the left-handed fermion doublets below, are column vectors. The Higgs field has mass dimension $1$, so the positive constant $M$ is a mass parameter. We will see later that the expansion parameter for BSM terms in our model is $\frac{v^{2}}{M^{2}}$, where $v$ is the Higgs vev, so we will need $M\gg v$.
@@
In our model, as in the Standard Model, it is this $U(2)$ group that is gauged.
@@
where the prefactor $M^{2}$ is required to obtain an expression of mass dimension $4$. This kinetic term has novel features compared with the Standard Model kinetic term $\overline{D_{\mu}\phi}D^{\mu}\phi$, but reduces to it if $|\phi|$ is small compared with $M$.
@@ 5 Higgs Potential
As usual, the constant, vacuum field configuration is selected to be $\phi=(\phi_{1},\phi_{2})=\left(0,\frac{1}{\sqrt{2}}v\right)$, which is charge-neutral and real. The unbroken gauge group is then the usual electromagnetic $U(1)_{\rm em}$ subgroup of $U(2)$ with typical element
@@ 7 Fermion Kinetic Terms
These assumptions mean that the kinetic terms (involving gauge-invariant spacetime derivatives) for the fermions are exactly as in the Standard Model, making use of the Dirac matrices $\gamma^{\mu}$ and $\gamma^{5}$. Fermion fields are doublet 4-spinors $\psi_{L}$ and singlet 4-spinors $\psi_{R}$ satisfying, respectively, $\gamma^{5}\psi_{L}=-\psi_{L}$ and $\gamma^{5}\psi_{R}=\psi_{R}$.
@@
For the first-generation lepton pair, the electron and its neutrino, the fields are the doublet $l_{L}=(\nu_{L},e_{L})$ and the one singlet $e_{R}$. For the second- and third-generation lepton pairs we again do not write the Lagrangians explicitly. The lepton kinetic terms are
@@ 8 Yukawa Terms
We now turn to our $\mathbb{CP}^{2}$ model. There is still no unification of $\phi$ and $\phi_{C}$, and two distinct types of Yukawa term are needed. We will use the homogeneous Higgs field $\Phi$ in our Yukawa terms, and must ensure that the result is gauge invariant and unchanged by rescaling by $\kappa$. The useful scale-invariant objects derived from $\Phi$ are the projectors
@@
Note that the projector $1-P$ acting on a right-handed fermion singlet produces a doublet and a singlet. The doublet plays the role in our $\mathbb{CP}^{2}$ model that the product of $\phi$ with the right-handed singlet does in the Standard Model, and its contraction with a left-handed fermion doublet is $U(2)$ invariant. One could generate further Yukawa terms by using the projected singlet, but these would be of higher order in $\phi$.
For the second type of Yukawa term, giving the u-quark its mass, we need to replace $\phi$ by its conjugate $\phi_{C}$, and use this to construct the conjugate projector $1-P_{C}$. The right-handed u-quark singlet is the 3-vector $(u_{R},0,0)$. Acting with $1-P_{C}$ on this, and contracting with $q_{L}$, gives the Yukawa term
@@ 9 Phenomenology Beyond the Standard Model
with its contributions defined in sections 5 – 8 above. Some of the apparent differences from the Standard Model can be removed by a recalibration of parameters and rescaling of the fields. But there remain some significant modifications to SM relations, and we will use these together with recent experimental results as reviewed by the Particle Data Group (PDG) [7] to place constraints on the BSM mass parameter $M$. Also, when the denominator factors are expanded out, there arise several dimension-6 terms that are Standard Model terms multiplied by $\bar{\phi}\phi$. In the following, we will explore the consequences of the $\mathbb{CP}^{2}$ SMEFT for (i) gauge boson masses, (ii) Higgs cubic and quartic couplings, (iii) Yukawa couplings. Constraints on $M$ currently come mainly from the $W$-boson to $Z$-boson mass ratio.
@@ 9.1 The $W$-boson to $Z$-boson mass ratio
The $W$-boson fields are $W_{\mu}^{\pm}=\frac{1}{\sqrt{2}}(A_{\mu}^{1}\mp A_{\mu}^{2})$. It is not surprising that both gauge-boson mass terms (49) have no contribution from the photon field. The photon remains massless because of the unbroken $U(1)_{\rm em}$ symmetry.
@@
$\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)^{2}}\biggl{[}\left(M^{2}+\frac{1}{2}v^{2}\right)\frac{v^{2}}{8}\left[g^{2}(A_{\mu}^{1})^{2}+g^{2}(A_{\mu}^{2})^{2}+(g^{2}+g^{\prime 2})Z_{\mu}^{2}\right]$
$\qquad\qquad\quad\qquad\qquad\qquad-\frac{v^{4}}{16}(g^{2}+g^{\prime 2})Z_{\mu}^{2}\biggr{]}.$ (52)
$v^{4}Z_{\mu}^{2}$ cancels here, and what remains can be reexpressed as
$\frac{1}{2}\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)^{2}}\left(\frac{gv}{2}\right)^{2}\left[\left(M^{2}+\frac{1}{2}v^{2}\right)((A_{\mu}^{1})^{2}+(A_{\mu}^{2})^{2})+\frac{M^{2}}{\cos^{2}\theta_{w}}Z_{\mu}^{2}\right].$ (53)
@@
From here on, we work to leading order in $\frac{v^{2}}{M^{2}}$, as our estimates below show that this quantity is less than $0.01$. The masses become
@@ 9.2 Higgs-boson mass and couplings
Here we consider the terms of the Lagrangian involving the Higgs field alone, leaving out couplings of the Higgs field to the gauge potentials and fermions.
@@
where $v_{\rm SM}$ is the SM Higgs vev and $H$ the real Higgs boson field. Expanding out, we obtain
@@
Unfortunately, the cubic and quartic Higgs couplings are not yet determined experimentally. There are only weak upper bounds on their strengths. In future, using the HL-LHC at CERN, it is hoped that these couplings can be measured and compared with SM predictions. Even then, the small modifications $\kappa_{3}$ and $\kappa_{4}$ to the SM predictions will probably remain out of reach. So these couplings give no useful constraint on $M$. However, the discussion here has clarified the need for the rescaling of the Higgs boson field $h$ as well as the Higgs vev $v$. These rescalings have an important effect on the Yukawa couplings, which we calculate next.
@@ 9.3 Higgs-fermion couplings
The kinetic terms for fermions have their canonical form in ${\mathcal{L}}$, so we need only investigate the Yukawa terms. To determine the fermion masses and coupling we need to retain only the terms up to linear order in the Higgs field $H$. As before, we convert the denominator factor in ${\mathcal{L}}_{\rm Yukawa}$ to a numerator factor and obtain for any fermion $f$,
@@
We next rescale $v$ and $h$ by the common factor $\left(1-\frac{v^{2}}{2M^{2}}\right)$ to obtain the normalised quantities $v_{\rm SM}$ and $H$. To linear order in $H$, the Yukawa term becomes

### v3

@@ 1 Introduction
The Lagrangian has terms involving: (i) Gauge-covariant derivatives of the Higgs field, (ii) a Higgs potential, (iii) Gauge field tensors, (iv) Dirac (kinetic) terms for fermions, (v) Yukawa terms coupling the Higgs field and fermions. We will discuss these in turn, and then explore some phenomenological consequences, in particular, for the $W$-boson to $Z$-boson mass ratio, for the couplings of the $W$- and $Z$-bosons to the physical Higgs boson [^2] The earlier published version, JHEP 04 (2025) 180, did not include this topic., and for the Yukawa couplings of the Higgs boson. These will lead to constraints on the BSM mass parameter $M$.
@@ 2 $\mathbb{CP}^{2}$ Geometry
The complex projective plane $\mathbb{CP}^{2}$ is obtained from flat complex 3-space, $\mathbb{C}^{3}$, by quotienting out the action of scaling by a complex number. Geometrically, $\mathbb{CP}^{2}$ is the manifold of complex lines [^3] We use the mathematician’s notion of complex line $\mathbb{C}$ to denote what is more commonly called the complex plane. through the origin in $\mathbb{C}^{3}$.
@@
Later, we will be using the projections of a generic vector $\Psi\in\mathbb{C}^{3}$, defined at location $Z\in\mathbb{C}^{3}$, on to the line through the origin and $Z$ and on to its complement, which is a copy of the tangent space to $\mathbb{CP}^{2}$. These projections are respectively
@@
Because these projections are unaffected by rescaling $Z$ by $\kappa$, they are defined at location $Z$ on $\mathbb{CP}^{2}$. More precisely, over $\mathbb{CP}^{2}$ there is its 1-dimensional tautological bundle, whose fibre at a point $Z\in\mathbb{CP}^{2}$ is the line in $\mathbb{C}^{3}$ through the origin and $Z$, and its 2-dimensional tangent bundle. The 3-dimensional, trivial $\mathbb{C}^{3}$ bundle over $\mathbb{CP}^{2}$ is the direct sum of these, so an element $\Psi$ in the trivial $\mathbb{C}^{3}$ bundle is the sum of its projections on to the tautological and tangent bundles.
@@ 3 Squashed 3-Spheres
whose angular part is the metric of the round 3-sphere of radius $r$. Each of these 3-spheres is a $U(2)$ orbit, and there is a unique $U(2)$ fixed point at $r=0$. Note that the ratio of the circle radius to the 2-sphere radius is a constant, independent of $r$.
@@
Here, the angular part is the metric of a squashed 3-sphere, a Hopf circle bundle over a round 2-sphere, where the ratio of the circle radius to the 2-sphere radius is $1/\sqrt{1+r^{2}}$, i.e. dependent on $r$. For us, the important observation is that for small positive $r$, the radius ratio is close to unity, so the squashed 3-sphere is approximately round. The consequence for our model is that the custodial $SO(4)$ symmetry – the symmetry of a round 3-sphere – will be approximately realised.
As $r\to\infty$, the radius ratio approaches zero, compatibly with the completion of $\mathbb{CP}^{2}$ at infinity being simply a round 2-sphere, which is a special 2-dimensional orbit of $U(2)$. This 2-sphere has radius $\frac{1}{2}$ and is at a finite distance from any other point in $\mathbb{CP}^{2}$.
@@ 4 Higgs Kinetic Term
where $\phi=(\phi_{1},\phi_{2})$ is a complex scalar doublet.444 All these multiplets, and the left-handed fermion doublets below, are column vectors. The Higgs field has mass dimension $1$, so the positive constant $M$ is a mass parameter. We will see later that the expansion parameter for BSM terms in our model is $\frac{v^{2}}{M^{2}}$, where $v$ is the Higgs vev, so we will need $M\gg v$.
@@
In our model, as in the SM, it is this $U(2)$ group that is gauged.
@@
where the prefactor $M^{2}$ is required to obtain an expression of mass dimension $4$. This kinetic term has novel features compared with the SM kinetic term $\overline{D_{\mu}\phi}D^{\mu}\phi$, but reduces to it if $|\phi|$ is small compared with $M$.
@@ 5 Higgs Potential
As usual, in unitary gauge, the vacuum field configuration is selected to be $\phi=(\phi_{1},\phi_{2})=\left(0,\frac{1}{\sqrt{2}}v\right)$, which is charge-neutral and real. The unbroken gauge group is then the electromagnetic $U(1)_{\rm em}$ subgroup of $U(2)$ with typical element
@@ 7 Fermion Kinetic Terms
These assumptions mean that the kinetic terms (involving gauge-covariant spacetime derivatives) for the fermions are exactly as in the Standard Model, making use of the Dirac matrices $\gamma^{\mu}$ and $\gamma^{5}$. Fermion fields are doublet 4-spinors $\psi_{L}$ and singlet 4-spinors $\psi_{R}$ satisfying, respectively, $\gamma^{5}\psi_{L}=-\psi_{L}$ and $\gamma^{5}\psi_{R}=\psi_{R}$.
@@
For the first-generation lepton pair, the electron and its neutrino, the fields are the doublet $l_{L}=(\nu_{L},e_{L})$ and the one singlet $e_{R}$. For the similar second- and third-generation lepton pairs we again do not write the Lagrangians explicitly. The lepton kinetic terms are
@@ 8 Yukawa Terms
We now turn to the $\mathbb{CP}^{2}$ model. There is still no unification of $\phi$ and $\phi_{C}$, and two distinct types of Yukawa term are needed. We will use the homogeneous Higgs field $\Phi$ in our Yukawa terms, and must ensure that the result is gauge invariant and unchanged by rescaling by $\kappa$. The useful scale-invariant objects derived from $\Phi$ are the projectors
@@
Note that the projector $1-P$ acting on a right-handed fermion singlet produces a doublet and a singlet. The doublet plays the role in the $\mathbb{CP}^{2}$ model that the product of $\phi$ with the right-handed singlet does in the Standard Model, and its contraction with the corresponding left-handed fermion doublet is $U(2)$ invariant. This is the term we use. One could generate further Yukawa terms by using the projected singlet, but these would be of higher order in $\phi$.
For the second type of Yukawa term, giving the u-quark its mass, we need to replace $\phi$ by its conjugate $\phi_{C}$, and use this to construct the conjugate projector $1-P_{C}$. The right-handed u-quark singlet is the 3-vector $(u_{R},0,0)$. Acting with $1-P_{C}$ on this, and again contracting with $q_{L}$, gives the Yukawa term
@@ 9 Phenomenology Beyond the Standard Model
with its contributions defined in sections 5 – 8 above. Some of the apparent differences from the Standard Model can be removed by a recalibration of parameters and rescaling of the fields. But there remain some significant modifications to SM relations, and we will use these together with recent experimental results as reviewed by the Particle Data Group (PDG) [7] to place constraints on the BSM mass parameter $M$. Also, when the denominator factors are expanded out, there arise several dimension-6 terms that are SM terms multiplied by $\bar{\phi}\phi$. In the following, we will explore the consequences of the $\mathbb{CP}^{2}$ SMEFT for (i) gauge boson masses, (ii) Higgs cubic and quartic couplings, (iii) Yukawa couplings, and (iv) Higgs boson couplings to gauge bosons. Constraints on $M$ currently come mainly from the $W$-boson to $Z$-boson mass ratio.
@@ 9.1 The $W$-boson to $Z$-boson mass ratio
The $W$-boson fields are $W_{\mu}^{\pm}=\frac{1}{\sqrt{2}}(A_{\mu}^{1}\mp iA_{\mu}^{2})$. It is not surprising that both gauge-boson mass terms (49) have no contribution from the photon field. The photon remains massless because of the unbroken $U(1)_{\rm em}$ symmetry.
@@
$\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)^{2}}\biggl{[}\left(M^{2}+\frac{1}{2}v^{2}\right)\frac{v^{2}}{8}\left[2g^{2}W_{\mu}^{-}W^{\mu+}+(g^{2}+g^{\prime 2})Z_{\mu}Z^{\mu}\right]$
$\qquad\qquad\quad\qquad\qquad\qquad-\frac{v^{4}}{16}(g^{2}+g^{\prime 2})Z_{\mu}Z^{\mu}\biggr{]}.$ (52)
$v^{4}Z_{\mu}Z^{\mu}$ cancels here, and what remains can be reexpressed as
$\frac{M^{2}}{\left(M^{2}+\frac{1}{2}v^{2}\right)^{2}}\left(\frac{gv}{2}\right)^{2}\left[\left(M^{2}+\frac{1}{2}v^{2}\right)W_{\mu}^{-}W^{\mu+}+\frac{1}{2}\frac{M^{2}}{\cos^{2}\theta_{w}}Z_{\mu}Z^{\mu}\right].$ (53)
@@
From here on, we work to leading non-trivial order in $\frac{v^{2}}{M^{2}}$, as the estimates below show that this quantity is less than $0.01$. The masses become
@@ 9.2 Higgs-boson mass and couplings
Here we consider the terms of the Lagrangian involving the Higgs field alone, leaving out couplings of the Higgs field to the gauge bosons and fermions.
@@
where $v_{\rm SM}$ is the SM Higgs vev and $H$ the physical Higgs boson field. Expanding out, we obtain
@@
Unfortunately, the cubic and quartic Higgs couplings are not yet determined experimentally. There are only weak upper bounds on their strengths. In future, using the HL-LHC at CERN, it is hoped that these couplings can be measured and compared with SM predictions. Even then, the small modifications $\kappa_{3}$ and $\kappa_{4}$ to the SM predictions will probably remain out of reach. So these couplings give no useful constraint on $M$. However, the discussion here has clarified the need for the rescaling of the Higgs boson field $h$ as well as the Higgs vev $v$. These rescalings have an important effect on the Yukawa couplings, which we calculate next, and on the couplings of the Higgs boson to the gauge bosons.
@@ 9.3 Higgs-fermion couplings
The kinetic terms for fermions have their canonical form in ${\mathcal{L}}$, so we need only investigate the Yukawa terms. To determine the fermion masses and couplings we need to retain only the terms up to linear order in the Higgs field $h$. As before, we convert the denominator factor in ${\mathcal{L}}_{\rm Yukawa}$ to a numerator factor and obtain for any fermion $f$,
@@
We next rescale $v$ and $h$ by the common factor $1-\frac{v^{2}}{2M^{2}}$ to obtain the canonical $v_{\rm SM}$ and $H$. To linear order in $H$, the Yukawa term becomes
@@
9.4 The $WWH$ and $ZZH$ couplings
Finally, we determine the tree-level couplings of the $Z$- and $W$-bosons to the physical Higgs boson. These couplings are found by expanding the Higgs kinetic Lagrangian (14) to linear order in the Higgs field $h$. The coupling of the photon to the Higgs boson vanishes at tree level. As before, we go to unitary gauge, writing $\phi=\frac{1}{\sqrt{2}}(0,v+h)$, and will rescale both $v$ and the field $h$ to obtain the canonical SM mass for the $Z$-boson and the canonical kinetic term for the Higgs field. The rescaled quantities are
$v_{\rm SM}=\left(1-\frac{v^{2}}{2M^{2}}\right)v\,,\quad H=\left(1-\frac{v^{2}}{2M^{2}}\right)h\,.$ (78)
The desired coupling terms are quadratic in the gauge boson fields and linear in the Higgs field, and arise from (14) after excluding the derivative $\partial_{\mu}h$. We set
$M^{2}+\bar{\phi}\phi=M^{2}+\frac{1}{2}(v+h)^{2}$ (79)
and from (14) read off the remaining part
$\frac{M^{2}}{\left(M^{2}+\frac{1}{2}(v+h)^{2}\right)^{2}}\times$
$\qquad\biggl{[}\left(M^{2}+\frac{1}{2}(v+h)^{2}\right)\frac{(v+h)^{2}}{8}\left[2g^{2}W_{\mu}^{-}W^{\mu+}+(g^{2}+g^{\prime 2})Z_{\mu}Z^{\mu}\right]$
$\qquad\qquad\qquad-\frac{(v+h)^{4}}{16}(g^{2}+g^{\prime 2})Z_{\mu}Z^{\mu}\biggr{]},$ (80)
where, similarly as before, the $(v+h)^{4}Z_{\mu}Z^{\mu}$ contributions cancel, leaving
$\frac{M^{2}}{\left(M^{2}+\frac{1}{2}(v+h)^{2}\right)^{2}}\biggl{[}M^{2}\frac{(v+h)^{2}}{8}\left[2g^{2}W_{\mu}^{-}W^{\mu+}+(g^{2}+g^{\prime 2})Z_{\mu}Z^{\mu}\right]$
$\qquad\qquad\quad\qquad\qquad\qquad+\frac{(v+h)^{4}}{8}g^{2}W_{\mu}^{-}W^{\mu+}\biggr{]}.$ (81)
Expanding in powers of $h$, we reproduce at zeroth order the masses of the gauge bosons, and at first order obtain the desired coupling terms
$\frac{1}{2}\frac{M^{4}}{(M^{2}+\frac{1}{2}v^{2})^{2}}\,g^{2}v\,W_{\mu}^{-}W^{\mu+}h+\frac{1}{4}\frac{M^{4}(M^{2}-\frac{1}{2}v^{2})}{(M^{2}+\frac{1}{2}v^{2})^{3}}\,(g^{2}+g^{\prime 2})v\,Z_{\mu}Z^{\mu}h\,.$ (82)
Now we convert to the rescaled quantities (78), and expand to linear order in $\frac{v_{\rm SM}^{2}}{M^{2}}$ to obtain
$\frac{1}{2}g^{2}v_{\rm SM}\,W_{\mu}^{-}W^{\mu+}H+\frac{1}{4}(g^{2}+g^{\prime 2})\left(1-\frac{v_{\rm SM}^{2}}{M^{2}}\right)v_{\rm SM}\,Z_{\mu}Z^{\mu}H\,.$ (83)
The couplings are therefore
$g_{WWH}=\frac{1}{2}g^{2}v_{\rm SM}\,,\quad g_{ZZH}=\frac{1}{4}(g^{2}+g^{\prime 2})\left(1-\frac{v_{\rm SM}^{2}}{M^{2}}\right)v_{\rm SM}\,.$ (84)
In the SM, these couplings are
$g_{WWH}=\frac{1}{2}g^{2}v_{\rm SM}\,,\quad g_{ZZH}=\frac{1}{4}(g^{2}+g^{\prime 2})v_{\rm SM}\,,$ (85)
so the $W$-boson coupling is unchanged at this order in the $\mathbb{CP}^{2}$ SMEFT, but the $Z$-boson coupling is reduced by the factor $1-\frac{v_{\rm SM}^{2}}{M^{2}}$.
The current experimental constraints on these couplings from ATLAS and CMS at the Large Hadron Collider are quite tight [7]. The measurements agree with the SM predictions to within 5%. If, for the sake of argument, one assumes the $Z$-boson coupling is less than 1% below the SM value, then $M$ would be constrained to be greater than 2.5 TeV.